OpenStudy (anonymous):

Which of the following shows that f(x) grows faster than g(x)? the limit as x goes to infinity of f(x)/g(x) = 1000 the limit as x goes to infinity of f(x)/g(x) = 0 the limit as x goes to infinity of f(x)/g(x) = infinity None of these

OpenStudy (solomonzelman):

your opinion?

OpenStudy (anonymous):

i dot quite recall the rule, i think it was something like if they equal a finite non zero number they grow at the same rate but i don't remember what infinity and zero ment

OpenStudy (anonymous):

so i know its not A , and probably not D

OpenStudy (solomonzelman):

lets think of this logically. if \(\large\displaystyle \lim_{ x~\rightarrow \infty}~(~f(x)~/~g(x)~)=0\) then g(x) is growing faster and that is why when you divide f(x) by g(x) you get zero because g(x) is being larger and larger....

OpenStudy (solomonzelman):

So it is definitely not B.

OpenStudy (solomonzelman):

right, or disagree ?

OpenStudy (anonymous):

i see your logic , i totally agree :D

OpenStudy (solomonzelman):

ok, good

OpenStudy (anonymous):

thanks for the help !

OpenStudy (solomonzelman):

are we done?

OpenStudy (anonymous):

yep , its clearly C

OpenStudy (solomonzelman):

Good!!

OpenStudy (solomonzelman):

\(\large\displaystyle \lim_{ x~\rightarrow \infty}~(~f(x)~/~g(x)~)=\infty\)

OpenStudy (solomonzelman):

yw

OpenStudy (anonymous):

:D

OpenStudy (solomonzelman):

Have you done LHospitals rule yet? just curious...